OpenStudy (anonymous):

Over the weekend, Claudia drove a total of 405 miles, driving twice as far on Saturday as on Sunday. How far did Claudia travel each day?

OpenStudy (anonymous):

If she drove x miles on Sunday, she drove twice as much (2x)on Saturday. So she drove a total of x+2x=405 miles 3x=405 x=405/3 x=135 miles